Selected data taken from the accounting records of Ginis Inc. for the current year ended December 31 are as follows:

Balance, December 31 Balance, January 1
Accrued expenses payable (operating expenses) $7,100 $7,760
Accounts payable (merchandise creditors) 38,910 42,950
Inventories 74,270 80,730
Prepaid expenses 4,000 4,800

During the current year, the cost of merchandise sold was $429,800, and the operating expenses other than depreciation were $96,000. The direct method is used for presenting the cash flows from operating activities on the statement of cash flows.

a. Determine the amount reported on the statement of cash flows for cash payments for merchandise.
$

b. Determine the amount reported on the statement of cash flows for cash payments for operating expenses.
$
Answer #1
a
Cost of merchandise sold 429800
Add: Decrease in Accounts payable 4040 =42950-38910
Less: Decrease in Inventories -6460 =74270-80730
Cash payments for merchandise 427380
b
Operating expenses 96000
Add: Decrease in Accrued expenses payable 660 =7760-7100
Less: Decrease in Prepaid expenses -800 =4000-4800
Cash payments for operating expenses 95860